Please help me solve 10x^2+3x=4
Answer by ankor@dixie-net.com(22740)

solve 10x^2+3x=4
subtract 4 from both sides
10x^2 + 3x - 4 = 0
you can use the quadratic formula a=10, b=3, c=-4. But this will factor to
(2x-1)(5x+4) = 0
Two solutions
2x = 1
x = 1/2
and
5x = -4
x = -4/5
:
you can check these solutions in the original problem